OpenStudy (popstar30977):

Simplify the expression. 24 x(1/4) ÷(-3) A.–18 B. –2 C. 2 D.18

OpenStudy (anonymous):

\[24\times \frac{ 1 }{ 4 }\div \left( -3 \right)=24\times \frac{ 1 }{ 4 }\times \frac{ 1 }{ -3 }=?\]

OpenStudy (popstar30977):



OpenStudy (popstar30977):

@surjithayer

OpenStudy (popstar30977):

@Freez @Brittnay @TheSmartOne @TheЅmartOne

OpenStudy (popstar30977):

@Austin6i6

OpenStudy (anonymous):

\[24\times \frac{ 1 }{ 4 }\times \frac{ 1 }{ -3 }\] \[\frac{24*1*1}{4*-3}\]

OpenStudy (popstar30977):

@jabez177 @maddiegirl @k_lynn

OpenStudy (popstar30977):

@Joky0424

OpenStudy (popstar30977):

@Joky0424

OpenStudy (anonymous):

\(\sf\Large\frac{24}{4 \times -3 }=?\)

OpenStudy (anonymous):

24x(1/4)/(-3) 6x/-3 -2x

OpenStudy (anonymous):

I am not sure about this though.